# Attention Sinks and Compression Valleys in LLMs are Two Sides of the Same Coin

### v1

@@ 3.1 Empirical Correlation: Synchronized Emergence Across Models
Figure 2: The coupled emergence of massive activations, compression, and sinks develops early in training and persists. Evolution of normalized entropy (left), sink rate (middle), and bos norm (right) across training checkpoints (1–143k steps) for Pythia 410M. All three phenomena emerge together around step 1k and remain synchronized throughout training, indicating this organization is learned early.
We observe that all three patterns align precisely. When the bos norm spikes to factors of $10^{3}$$10^{4}$ (typically layers 0–5 depending on model depth), entropy simultaneously drops and sink rates surge. We compute the Pearson correlation between the change in bos norm and entropy, obtaining $r=-0.9\pm 0.18$ across models, while bos norm and sink rate correlate at $r=0.58\pm 0.25$.
@@ 3.2 Theoretical Framework: Massive Activations Imply Compression
By the characterization of singular values, $\sigma_{1}^{2}=\max_{\|\mathbf{v}\|=1}\|\mathbf{X}\mathbf{v}\|^{2}$. Choosing $\mathbf{v}=\mathbf{x}_{0}/\|\mathbf{x}_{0}\|$ and expanding $\|\mathbf{X}\mathbf{v}\|^{2}$ yields the bound. Full proofs and discussions in Appendix A.2. ∎
@@
Figure 3: Our theoretical bounds become exact when massive activations emerge, proving they drive compression. Left: When bos norm dominates (layers 5–15), the first singular value $\sigma_{1}^{2}$ approximately equals both $\|{\mathbf{x}}_{\text{BOS}}\|^{2}$ and $\|{\mathbf{X}}\|_{F}^{2}$, confirming near rank-one structure. Right: Our entropy upper bound (Corollary 2) tightly matches empirical values in compressed layers, validating that massive activations mathematically necessitate compression. Average across 100 RedPajama (Weber et al., 2024) examples for Pythia 410M.
@@ 3.3 Evidence from Targeted Ablations
We find that ablating massive activations can eliminate both phenomena, confirming our theoretical conjecture. In LLaMA3 8B, removing the MLP’s contribution at the first layer prevents entropy drop (remains at 0.4-0.5 bits vs. dropping to 0.02 bits), eliminates sink formation (sink rate drops from 0.85-1.0 to 0.0) and maintains bos norm within $2\times$ of other tokens (vs. $10^{2}\times$ normally) (Fig. 4). Similar results hold across models, as seen in Fig.14 in Appendix B.1. Some models (Pythia 410M, Qwen2 7B) develop massive activations in more than one stage. Ablating any single stage partially reduces compression; ablating all stages eliminates it entirely, suggesting a cumulative contribution. However, for Pythia 410M, ablations remove compression but do not remove attention sinks, which suggests that the formation of sinks might have model-dependent causes.
@@ 4.3 Phase 3: Selective Refinement (Layers 85–100%)
Norm equalization drives decompression. In this phase, we find the bos norm plateaus or decreases while the average norm of the rest of tokens rises sharply, driving them toward similar magnitudes (right panel in Figure 4 and Figure 15 in Appendix B.1). This equalization begins earlier than the full phase transition, where the average norm starts rising around 40–60% depth, preparing for the eventual shift. The massive activation ratio drops from $10^{3}$ to <10<10, removing the mathematical basis for compression and allowing representations to re-expand.
@@ 5.1 The Distinct Performance Patterns Across Tasks
Embedding tasks peak in middle layers. To highlight the difference between generation and embedding tasks, we implement the following linear probing experiment. For each dataset, we encode the examples with the frozen backbone and extract hidden states at every layer. We then train a linear classifier at each layer on the training split and evaluate it on the test split. We probe the same multiple-choice QA benchmarks as before and, additionally, a sentence classification dataset (SST-2; Socher et al. 2013), assessing how much task-relevant information is linearly accessible at different depths. Figure 7 (right) shows the results for ARC Easy, highlighting the difference in optimal depths with the corresponding generation task, while full results for this experiment are shown in Fig. 26, Appendix B.3. Moreover, we reproduce Skean et al. (2025) on 32 MTEB tasks (Muennighoff et al., 2022) across a broader set of larger, decoder-only models. Across the board, we find that performance peaks consistently at 25–75% relative depth, outperforming early/late layers by 10-20% and precisely aligning with Phase 2, where compression is strongest (see Fig. 27 in Appendix B.3). These results align with evidence that next-token pretraining does not uniformly benefit perception-style classification (Balestriero & Huang, 2024). Together, they underscore task dependence: classification-relevant linear features concentrate in intermediate layers, whereas late layers are repurposed for token-specific generative refinement. Furthermore, the pattern suggests that massive activations in the residual pathway not only curb over-mixing via sink formation but also act as a mechanism the model uses to compress information.
Figure 7: Embedding tasks peak during compression while generation requires full refinement, revealing distinct computational objectives. (Left, Middle) Perplexity on Wikitext-2 and multiple-choice QA accuracy in ARC Easy via LogitLens generally do not improve significantly until $\sim$50% depth, then decreases/rises steadily through Phase 3. (Right) Linear probe test accuracy on the same task peaks at 25–75% depth (Phase 2) and declines thereafter. This divergence demonstrates that embedding-relevant features concentrate in compressed middle layers, while generation tasks require full-depth for token-specific predictions.
@@ 5.2 Why Do Different Tasks Need Different Phases?
We believe these findings clarify which tasks actually benefit from compression. In particular, embedding-style objectives (such as clustering, retrieval, classification, bitext mining, etc.) gain from Phase 2’s compression because they target low-dimensional structure while discarding irrelevant information, echoing classic arguments on the benefits of information bottlenecks and compressed representations (Shwartz-Ziv et al., 2018; Kawaguchi et al., 2023). This picture aligns with evidence that LLMs produce surprisingly linear (and often linearly separable) embeddings (Razzhigaev et al., 2024; Marks & Tegmark, 2023). In particular, when features concentrate in a low-dimensional subspace, linear probing, semantic retrieval, and related embedding tasks become easier. Moreover, such linear structure has been linked to the emergence of high-level semantic concepts (Park et al., 2023), reinforcing our hypothesis on why mid-layer compressed states tend to work well for non-generative evaluations.
@@ Appendix B Additional results
All experiments were implemented in PyTorch using NVIDIA A100 GPUs with 40GB memory or NVIDIA H100 GPUs with 80GB when the memory requirements were stronger. We examined pretrained models of varying depths, using HuggingFace repositories with Transformers and Transformer-Lens (Nanda & Bloom, 2022). When large datasets were run to collect metrics such as sink rates and norms, prompts were truncated to a maximum length of 4096 tokens for the FineWeb-Edu experiment (Fig. 5) and 1024 for the GSM8K experiment (Fig. 1), as the latter required singular value decompositions to compute the entropy. LogitLens experiments for multiple-choice-questions tasks were done with LM-Evaluation-Harness (Gao et al., 2024), implementing our own model wrapper to output hidden states at each layer instead of final ones.
@@ B.1 Broader Analysis of Models
Validation on more model families and large models.
@@
We evaluated the training dynamics of the Pythia 410M/6.9B/12B models across multiple checkpoints (steps 1, 1k, 2k, 4k, 8k, 10k, 20k, 30k, and 143k). At each checkpoint, after every layer we recorded the entropy, bos sink rate (threshold $\tau=0.3$) and the norm of the bos token representation. The prompt was a single GSM8K prompt “Janet’s ducks lay 16 eggs...” Figures 10 and 11 illustrate the results for the Pythia 6.9B and 12B models, respectively.
Figure 10: Entropy, Sink rate and bos Norm at different training checkpoints for Pythia 6.9B.
@@
Visualizations of theoretical results.
@@
Figure 12: Theoretical bounds for Pythia 410M.
Figure 13: Theoretical bounds for LLaMA3 8B.
@@
We further run the targeted MLP ablations on more models to erase the appearance of the massive activation. For LLaMA3 8B, we ablate layer 0; for Qwen2 7B, we ablate layers 3 and 4 and for Pythia 410M, we ablate layers 5-7. The results are shown in figure 14. Interestingly, removing massive activation always decompresses the representations, however, in Pythia 410M it does not remove the attention sinks, which might be explained by the many architectural differences between these models.
Figure 14: MLP ablations for LLaMA3 8B, Qwen2 7B and Pythia 410M.
Norm equalizations.
Across models, we find that the average norm of the rest of the tokens (excluding bos) grows monotonically with depth, while the bos norm grows abruptly with the massive activation, remains constant in the middle layers and drops at the last layers. Figure 15 illustrates this process for three models. As the rest of tokens become closer to the bos norm, the dominance of the first singular value is weaker, allowing for representations to decompress.
Figure 15: LLMs equalize token norms towards at the end of the network.
@@ B.2 Mixing and Sink Detection Metrics
Figures 21 and 22 show the FineWeb-Edu experiment for Bloom 1.7B and Qwen2 7B models. The trend is clear: regardless of the input, the models do not allocate attention to the bos token until the massive activation emerges. The amount of sinks present in the middle layers is input-dependent, however the amount of mixing performed in the early layers is not.
@@ B.3 Additional results on downstream tasks
To further validate the results of Section 5 and the ones proposed by (Skean et al., 2025), we run a standard linear probing experiment. Probes are trained independently per layer, with backbone parameters fixed, using a learning rate of $5\times 10^{-4}$, one epoch, maximum length of 1024 and batch sizes of 16-32. We train for backbones Pythia 410M, Pythia 6.9B, LLaMA3 8B, Qwen2 7B and Gemma 7B. Figure 26 shows the results. As discussed, across models and datasets, accuracy peaks in the middle layers. These results suggest that the linear features relevant for classification emerge transiently in the compressed middle representations, while the late layers are repurposed for generative refinement. Moreover, we run 32 MTEB tasks for the same models and report the average main score across tasks in Figure 27.

### v2

@@ 3.1 Empirical Correlation: Synchronized Emergence Across Models
Figure 2: The coupled emergence of massive activations, compression, and sinks develops early in training and persists. Evolution of normalized entropy (left), sink rate (middle), and bos norm (right) across training checkpoints (1–143k steps) for Pythia 410M, 6.9B. All three phenomena emerge together around step 1k and remain synchronized throughout training, indicating this organization is learned early.
We observe that all three patterns align precisely. When the bos norm spikes to factors of $10^{3}$$10^{4}$ (typically layers 0–5 depending on model depth), entropy simultaneously drops and sink rates surge. We compute the Pearson correlation between the change in bos norm and entropy, obtaining $r=-0.9\pm 0.18$ across models, while bos norm and sink rate correlate at $r=0.58\pm 0.25$ [^1] Here, Llama-3-8B behaves as an extreme outlier: almost all layers exhibit very large BoS activations, which makes the layerwise first-order differences nearly flat. This, in turn, renders the per-layer correlation ill-defined and artificially suppresses the overall correlation magnitude while inflating its variance..
@@ 3.2 Theoretical Framework: Massive Activations Imply Compression
Full proofs and discussions in Appendix A.2. ∎
@@
Figure 3: Our theoretical bounds become exact when massive activations emerge, proving they drive compression. Left: When bos norm dominates (layers 5–15), the first singular value $\sigma_{1}2$ approximately equals both $\|{\mathbf{x}}_{\text{BOS}}\|$ and $\|{\mathbf{X}}\|_{F}$, confirming near rank-one structure. Right: Our entropy upper bound (Corollary 2) tightly matches empirical values in compressed layers, validating that massive activations mathematically necessitate compression. Average across 100 RedPajama (Weber et al., 2024) examples for Pythia 410M.
@@ 3.3 Evidence from Targeted Ablations
We find that ablating massive activations can eliminate both phenomena, confirming our theoretical conjecture. In LLaMA3 8B, removing the MLP’s contribution at the first layer prevents entropy drop (remains at 0.4-0.5 bits vs. dropping to 0.02 bits), eliminates sink formation (sink rate drops from 0.85-1.0 to 0.0) and maintains bos norm within $2\times$ of other tokens (vs. $10^{2}\times$ normally) (Fig. 4). Similar results hold across models, as seen in Figs.17,18 in Appendix B.1. Some models (Pythia 410M, Qwen2 7B) develop massive activations in more than one stage. Ablating any single stage partially reduces compression; ablating all stages eliminates it entirely, suggesting a cumulative contribution.
@@ 4.3 Phase 3: Selective Refinement (Layers 85–100%)
Norm equalization drives decompression. In this phase, we find the bos norm plateaus or decreases while the average norm of the rest of tokens rises sharply, driving them toward similar magnitudes (right panel in Figure 4 and Figure 7). This equalization begins earlier than the full phase transition, where the average norm starts rising around 40–60% depth, preparing for the eventual shift. The massive activation ratio drops from $10^{3}$ to <10<10, removing the mathematical basis for compression and allowing representations to re-expand.
@@
Figure 7: LLMs equalize token norms towards at the end of the network.
@@ 5.1 The Distinct Performance Patterns Across Tasks
Embedding tasks peak in middle layers. To highlight the difference between generation and embedding tasks, we implement the following linear probing experiment. For each dataset, we encode the examples with the frozen backbone and extract hidden states at every layer. We then train a linear classifier at each layer on the training split and evaluate it on the test split. We probe the same multiple-choice QA benchmarks as before and, additionally, a sentence classification dataset (SST-2; Socher et al. 2013), assessing how much task-relevant information is linearly accessible at different depths. Figure 8 (right) shows the results for ARC Easy, highlighting the difference in optimal depths with the corresponding generation task, while full results for this experiment are shown in Fig. 31, Appendix B.3. Moreover, we reproduce Skean et al. (2025) on 32 MTEB tasks (Muennighoff et al., 2022) across a broader set of larger, decoder-only models. Across the board, we find that performance peaks consistently at 25–75% relative depth, outperforming early/late layers by 10-20% and precisely aligning with Phase 2, where compression is strongest (see Fig. 32 in Appendix B.3). These results align with evidence that next-token pretraining does not uniformly benefit perception-style classification (Balestriero and Huang, 2024). Together, they underscore task dependence: classification-relevant linear features concentrate in intermediate layers, whereas late layers are repurposed for token-specific generative refinement. Furthermore, the pattern suggests that massive activations in the residual pathway not only curb over-mixing via sink formation but also act as a mechanism the model uses to compress information.
Figure 8: Embedding tasks peak during compression while generation requires full refinement, revealing distinct computational objectives. (Left, Middle) Perplexity on Wikitext-2 and multiple-choice QA accuracy in ARC Easy via LogitLens generally do not improve significantly until $\sim$50% depth, then decreases/rises steadily through Phase 3.
(Right) Linear probe test accuracy on the same task peaks at 25–75% depth (Phase 2) and declines thereafter. This divergence demonstrates that embedding-relevant features concentrate in compressed middle layers, while generation tasks require full-depth for token-specific predictions.
@@ 5.2 Why Do Different Tasks Need Different Phases?
We believe these findings clarify which tasks actually benefit from compression. In particular, embedding-style objectives (such as clustering, retrieval, classification, bitext mining, etc.) gain from Phase 2’s compression because they target low-dimensional structure while discarding irrelevant information, echoing classic arguments on the benefits of information bottlenecks and compressed representations (Shwartz-Ziv et al., 2018; Kawaguchi et al., 2023). This picture aligns with evidence that LLMs produce surprisingly linear (and often linearly separable) embeddings (Razzhigaev et al., 2024; Marks and Tegmark, 2023). In particular, when features concentrate in a low-dimensional subspace, linear probing, semantic retrieval, and related embedding tasks become easier. Moreover, such linear structure has been linked to the emergence of high-level semantic concepts (Park et al., 2023), reinforcing our hypothesis on why mid-layer compressed states tend to work well for non-generative evaluations.
@@ Appendix B Additional results
All experiments were implemented in PyTorch using NVIDIA A100 GPUs with 40GB memory or NVIDIA H100 GPUs with 80GB when the memory requirements were stronger. We examined pretrained models of varying depths, using HuggingFace repositories with Transformers and Transformer-Lens (Nanda and Bloom, 2022). When large datasets were run to collect metrics such as sink rates and norms, prompts were truncated to a maximum length of 4096 tokens for the FineWeb-Edu experiment (Fig. 5) and 1024 for the GSM8K experiment (Fig. 1), as the latter required singular value decompositions to compute the entropy. LogitLens experiments for multiple-choice-questions tasks were done with LM-Evaluation-Harness (Gao et al., 2024), implementing our own model wrapper to output hidden states at each layer instead of final ones.
@@ B.1 Broader Analysis of Models
Validation on more datasets.
The GSM8K dataset use in Figure 1 only contains structured question-answer pairs that we concatenate to obtain each prompt. In order to test our theory against different distributions of data, we further evaluate these models on 5K examples from the following datasets: The Stack (Kocetkov et al., 2022), for code; FineWeb (Penedo et al., 2024), for prose and the MATH dataset (Hendrycks et al., 2021), for math questions. The results are shown in Figures 9,10 and 11. We find that the results to be very similar to those of GSM8K, further reinforcing the input-independence of our theory.
Figure 9: Entropy, sink rate and bos in The Stack.
Figure 10: Entropy, sink rate and bos in FineWeb.
Figure 11: Entropy, sink rate and bos in the MATH dataset.
Validation on more model families and large models.
@@
We evaluated the training dynamics of the Pythia 410M/6.9B/12B models across multiple checkpoints (steps 1, 1k, 2k, 4k, 8k, 10k, 20k, 30k, and 143k). At each checkpoint, after every layer we recorded the entropy, bos sink rate (threshold $\tau=0.3$) and the norm of the bos token representation. The prompt was a single GSM8K prompt “Janet’s ducks lay 16 eggs...” Figure 14 illustrates the results for the Pythia 12B model.
@@
Role of $\alpha$ in practice and visualization of theoretical results.
@@
Figure 15: Theoretical bounds for Pythia 410M. (a) shows the elements in the bound of Theorem 1. (b) shows the singular value spectrum of $\mathbf{X}^{(\ell)}$ at each layer $\ell$. For plots (c,d,e) gray represents the weaker bound (no $\alpha$), while green represents the bound as presented in Corollary 2.
Figure 16: Theoretical bounds for LLaMA3 8B. (a) shows the elements in the bound of Theorem 1. (b) shows the singular value spectrum of $\mathbf{X}^{(\ell)}$ at each layer $\ell$. For plots (c,d,e) gray represents the weaker bound (no $\alpha$), while green represents the bound as presented in Corollary 2.
@@
We further run the targeted MLP ablations on more models to erase the appearance of the massive activation. For LLaMA3 8B, we ablate layer 0; for Qwen2 7B, we ablate layers 3 and 4 and for Pythia 410M, we ablate layers 0 and 5-7. The results are shown in Figures 4, 17, 18.
Figure 17: MLP ablation for Qwen2 7B.
Figure 18: MLP ablations for Pythia 410M.
Additional Ablations on Massive Activations
Here, we also run ablation regarding the emergence and exact effect of massive activation on both compression and attention sinks on Pythia 410M. In particular, Figure 19 shows what components of the model cause the emergence of a massive activation on the BoS token. On the other hand, Figure 20 shows the effect of clipping the maximum value in the massive activation vector to be between $\gamma,\gamma]$.
Figure 19: MLPs are responsible for the appearance of massive activations in Pythia 410M.
Figure 20: Ablation of the effects of massive activations in the residual pathway on Pythia 410M. Lower massive activation norms generally correspond to both increased compression and lower sinks.
From the Figures, one can see that the MLP seems to be causing the emergence of massive activations and how lowering the norm of the massive activation corresponds to an increase the in the representational entropy and decrease in attention sinks. In general, we find attention sinks to be less predictable than entropy (perhaps due to the way the sink score is calculated), although there is a general downward trend as the massive activation norm is reduced.
@@ B.2 Mixing and Sink Detection Metrics
Figures 26 and 27 show the FineWeb-Edu experiment for Bloom 1.7B and Qwen2 7B models. The trend is clear: regardless of the input, the models do not allocate attention to the bos token until the massive activation emerges. The amount of sinks present in the middle layers is input-dependent, however the amount of mixing performed in the early layers is not.
@@ B.3 Additional results on downstream tasks
To further validate the results of Section 5 and the ones proposed by (Skean et al., 2025), we run a standard linear probing experiment. Probes are trained independently per layer, with backbone parameters fixed, using a learning rate of $5\times 10^{-4}$ with PyTorch’s default settings for Adam (Kingma and Ba, 2015), no weight decay, one epoch, maximum length of 1024, batch sizes of 16-32 and a random fixed seed of 123. We train for backbones Pythia 410M, Pythia 6.9B, LLaMA3 8B, Qwen2 7B and Gemma 7B. Figure 31 shows the results. As discussed, across models and datasets, accuracy peaks in the middle layers. These results suggest that the linear features relevant for classification emerge transiently in the compressed middle representations, while the late layers are repurposed for generative refinement. Moreover, we run 32 MTEB tasks for the same models and report the average main score across tasks in Figure 32.